# Forward and Backward State Abstractions for Off-policy Evaluation

### v1

@@ D.3 Proof of Theorem 3
    where the first equation is due to $\rho^{\pi}$-irrelevance and the second equation follows from (8). This yields
    $\rho^{\pi}(A_{t-1},S_{t-1})\perp\!\!\!\perp S_{t}|\phi(S_{t}).$
    We assume that for $k\leq t-2$ the formulation (D.8) holds. Now, we prove that for $k=t-1$, (D.8) successes. By similar arguments to that of (D.3), we get
    $\mathbb{P}\big{(}\rho^{\pi}(A_{1},S_{1})=c|S_{t}\big{)}=$ $\mathbb{P}[\mathbb{P}\big{(}\rho^{\pi}(A_{1},S_{1})=c|S_{2},A_{2},S_{t},A_{t}\big{)}|S_{t}]$
    $=$ $\mathbb{P}[\mathbb{P}\big{(}\rho^{\pi}(A_{1},S_{1})=c|S_{2}\big{)}|S_{t}]$
    $=$ $\mathbb{P}[g\big{(}\phi(S_{2})\big{)}|S_{t}].$ (D.10)
    To prove this, we need to show that for any $11\leq k\leq t-$ 1, we have
    $\mathbb{P}\big{(}\phi(S_{t-k})=x|S_{t}\big{)}=\mathbb{P}\big{(}\phi(S_{t-k})=x|\phi(S_{t})\big{)}.$ (D.11)
    The definition of inverse model implies when $k=1$, (D.11) successes. We assume that for $k\leq t-2$ the formulation (D.11) successes. Now, we prove that for $k=t-1$, (D.11) also hold.
    $\mathbb{P}\big{(}\phi(S_{1})=x|S_{t}\big{)}=$ $\mathbb{P}[\mathbb{P}\big{(}\phi(S_{1})=x|S_{2},S_{t}\big{)}|S_{t}]$
    $=$ $\underbrace{\mathbb{P}[\mathbb{P}\big{(}\phi(S_{1})=x|S_{2}\big{)}|S_{t}]}_{\mbox{Inverse Markovianity}}$
[…]
    $=$ $\mathbb{P}[g\big{(}\phi(S_{2})\big{)}|\phi(S_{t})].$
    This proves (D.11). Combing (D.3) and (D.11), we can get
    $\mathbb{P}\big{(}\rho^{\pi}(A_{1},S_{1})=c|S_{t}\big{)}=$ $\mathbb{P}[g\big{(}\phi(S_{2})\big{)}|\phi(S_{t})].$
    Then we prove (D.7).
    (2)In the second step, we will prove that if $\phi$ satisfies equation (D.7) and $\rho^{\pi}$-irrelevance, it is $w^{\pi}$-irrelevant, namely for any $s^{(1)}$ and $s^{(2)}$ satisfying $\rho_{t}^{\pi}(a,s^{(1)})=\rho_{t}^{\pi}(a,s^{(2)})$, they will satisfy
    $w^{\pi}(a,s^{(1)})=w^{\pi}(a,s^{(2)}).$
    It follows from the definition of state abstraction, $s^{(1)}$ and $s^{(2)}$, we have
    $\mathbb{P}(X_{t}|S_{t}=s^{(1)})={\bf{1}}\big{(}s^{(1)}\in\phi^{-1}(X_{t})\big{)}={\bf{1}}\big{(}s^{(2)}\in\phi^{-1}(X_{t})\big{)}=\mathbb{P}(X_{t}|S_{t}=s^{(2)}).$ (D.12)
    By (D.12) and (D.7), we have
    $w^{\pi}(a,s^{(1)})=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\mathbb{P}^{\pi}(A_{t}=a,S_{t}=s^{(1)})}{\mathbb{P}(A=a,S=s^{(1)})}$
    $=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\mathbb{P}^{\pi}(A_{t}=a|S_{t}=s^{(1)})\mathbb{P}^{\pi}(S_{t}=s^{(1)})}{\mathbb{P}(A=a|S=s^{(1)})\mathbb{P}^{b}(S=s^{(1)})}$
    $=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\rho_{t}^{\pi}(a,s^{(1)})\mathbb{P}^{\pi}(S_{t}=s^{(1)})}{\mathbb{P}^{b}(S=s^{(1)})}$
    $=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\rho_{t}^{\pi}(a,s^{(1)})\mathbb{E}^{\pi}[{\bf{1}}(S_{t}=s^{(1)})]}{\mathbb{E}^{b}[{\bf{1}}(S_{t}=s^{1})]}$
    $=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\rho_{t}^{\pi}(a,s^{(1)})\mathbb{E}^{b}[{\bf{1}}(S_{t}=s^{(1)})\prod_{j=1}^{t-1}\rho_{j}^{\pi}(A_{j},S_{j})]}{\mathbb{E}^{b}[{\bf{1}}(S_{t}=s^{(1)})]}$
    $=$ $\frac{(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\rho_{t}^{\pi}(a,s^{(1)})\mathbb{E}^{b}\left[\mathbb{E}^{b}\left({\bf{1}}(S_{t}=s^{(1)})\prod_{j=1}^{t-1}\rho_{j}^{\pi}(A_{j},S_{j})|X_{t}\right)\right]}{\mathbb{E}^{b}[{\bf{1}}(S_{t}=s^{(1)})]}$
[…]
    $=$ $(1-\gamma)\sum_{t=1}^{T}\gamma^{t-1}\rho_{t}^{\pi}(a,s^{(1)})\mathbb{E}^{b}\left(\frac{\mathbb{P}(X_{t}|S_{t}=s^{(1)})\prod_{j=1}^{t-1}\rho_{j}^{\pi}(A_{j},S_{j})}{\mathbb{P}(X_{t})}\right)$
[…]
    $=$ $w^{\pi}(a,s^{(2)}).$
    Then, we can conclude that backward-model-irrelevance implies the $\rho^{\pi}$-irrelevance and $w^{\pi}$-irrelevance.
  • It follows from the definition of $Q$-function-based method that
    $\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]=$ $\sum_{a,x}Q_{\phi}^{\pi}(a,x)\pi(a|x)\mathbb{P}(\phi(S_{1})=x)$
    $=$ $\sum_{a,x}\mathbb{E}^{\pi}\Big{[}\sum_{t=1}^{+\infty}\gamma^{t-1}R_{t}|X_{1}=x,A_{1}=a\Big{]}\pi(a|x)\mathbb{P}(X_{1}=x)$
    $=$ $\sum_{a,x,r}\sum_{t=1}^{+\infty}\gamma^{t-1}r\mathbb{P}^{\pi}\Big{[}r|X_{1}=x,A_{1}=a\Big{]}\pi(a|x)\mathbb{P}(X_{1}=x)$
    $=$ $\mathbb{E}^{\pi}\Big{[}\sum_{t=1}^{+\infty}\gamma^{t-1}R_{t}\Big{]}$
    $=$ $\mathbb{E}[f_{1}(Q^{\pi})].$
@@
  • The conclusion directly follows from the last conclusion of Theorem 1, and the first conclusion of Theorem 3.

### v2

@@ D.3 Proof of Theorem 1
    where the second last equation follows from $w_{t}^{\pi}$-irrelevance, $\rho^{\pi}$-irrelevance and (D.10). This yields $w_{t+1}^{\pi}$-irrelevance, and subsequently $w^{\pi}$-irrelevance, by induction. By Lemma D.1, $w^{\pi}$-irrelevance further yields $\mathbb{E}[f_{3}(w^{\pi})]=\mathbb{E}[f_{3}(w^{\pi}_{\phi})]$.
    It remains to prove the identifiability of $w_{\phi}^{\pi}$. We similarly define
    $w_{t,\phi}^{\pi}(a,x)=\frac{\mathbb{P}^{\pi}(A_{t}=a,\phi(S_{t})=x)}{\mathbb{P}(A_{t}=a,\phi(S_{t})=x)}.$
    It follows that $w_{\phi}^{\pi}=\sum_{t\geq 1}\gamma^{t-1}w_{t,\phi}^{\pi}$. Again, $w_{1,\phi}^{\pi}$ corresponds to $\rho_{\phi}^{\pi}$, which is identifiable under backward-model-irrelevance and $\pi$-irrelevance.
    Based on the aforementioned arguments, we can show that
    $w_{t+1,\phi}^{\pi}(a,x)=\sum_{a^{\prime},x^{\prime}}w_{t,\phi}^{\pi}(a^{\prime},x^{\prime})\rho_{\phi}(a|x)\mathbb{P}(A_{t}=a^{\prime},\phi(S_{t})=x^{\prime}|A_{t+1}=a,\phi(S_{t+1})=x),$
    where the last term on the RHS is well-defined according to (D.10). Suppose we have shown the identifiability of $w_{t,\phi}^{\pi}$. Then each term on the RHS is identifiable. This proves the identifiability of $w_{t+1,\phi}^{\pi}$. By induction, $w_{t,\phi}^{\pi}$ is identifiable for each $t\geq 1$.
    According to the coverage and stationarity assumptions in Assumptions 2 and 3, the denominators in $\{w_{t,\phi}^{\pi}\}$ are bounded away from zero. Consequently, $\{w_{t,\phi}^{\pi}\}$ are uniformly bounded. By letting $t\to\infty$, we obtain the identifiability of $w_{\phi}^{\pi}$. The proof is thus completed.
  • Fisher consistency of Q-function-based method. We first show that $\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]=\mathbb{E}[f_{1}(Q^{\pi})]$ under $\pi$-irrelevance. This is immediate by noting that
    $\mathbb{E}[f_{1}(Q^{\pi})]=J(\pi)=\sum_{t\geq 1}\gamma^{t-1}\mathbb{E}^{\pi}(R_{t})=\sum_{a,x}\sum_{t\geq 1}\gamma^{t-1}\mathbb{E}^{\pi}(R_{t}|A_{1}=a,\phi(S_{1})=x)$
    ×$\times\mathbb{P}^{\pi}(A_{1}=a|\phi(S_{1})=x)\mathbb{P}(\phi(S_{1})=x)=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})],$
    where the first term on the second line equals $\pi(a|s)$ for any $s$ such that $\phi(s)=x$, under $\pi$-irrelevance.
    It remains to prove the identifiability of $Q_{\phi}^{\pi}$ under $\pi$- and backward-model-irrelevance. First, we establish the identifiability of $\mathbb{E}^{\pi}(R_{1}|A_{1}=a,\phi(S_{1})=x)$. By definition
    $\mathbb{E}^{\pi}(R_{1}|A_{1}=a,\phi(S_{1})=x)=\frac{\mathbb{E}^{\pi}[R_{1}\mathbb{I}(A_{1}=a,\phi(S_{1})=x)]}{\mathbb{P}^{\pi}(A_{1}=a,\phi(S_{1})=x)}.$
    Using the change of measure theorem, the numerator equals
    $\mathbb{E}\Big{[}\rho^{\pi}(a|S_{1})R_{1}\mathbb{I}(A_{1}=a,\phi(S_{1})=x)\Big{]}=\mathbb{E}\Big{[}\rho_{\phi}^{\pi}(a|x)R_{1}\mathbb{I}(A_{1}=a,\phi(S_{1})=x)\Big{]}$
    $=\rho_{\phi}^{\pi}(a|x)\mathbb{E}\Big{[}R_{1}\mathbb{I}(A_{1}=a,\phi(S_{1})=x)\Big{]},$
    where the first equation holds due to $\pi$- and behavior-policy-irrelevance. Notice that the denominator equals $\mathbb{P}(\phi(S_{1})=x)\pi_{\phi}(a|x)$, it follows that
    $\mathbb{E}^{\pi}(R_{1}|A_{1}=a,\phi(S_{1})=x)=\mathbb{E}(R_{1}|A_{1}=a,\phi(S_{1})=x),$
    which is identiable from the data.
    Similarly, one can show that $\mathbb{P}^{\pi}(\phi(S_{2})=x^{\prime}|A_{1}=a,\phi(S_{1})=x)=\mathbb{P}(\phi(S_{2})=x^{\prime}|A_{1}=a,\phi(S_{1})=x)$ is identifiable as well.
    Now, the identifiability can be readily obtained if we show $(\phi(S_{t}),A_{t},R_{t})_{t\geq 1}$ remains an MDP. In that case, standard Q-learning algorithms can be applied to such a reduced MDP to consistently identify $Q_{\phi}^{\pi}$. Such an MDP property will be proven in Section D.4.2 under a more challenging setting that allows the behavior policy to be history-dependent.
@@
  • Fisher consistency of DRL. Due to the double robustness property of DRL, the conclusion directly follows from the last conclusion of Theorem D.1 and the first two conclusions of Theorem 1.